OpenStudy (anonymous):

Find the inverse function of f(x) = 7x – 9

OpenStudy (anonymous):

just like solving \[7x-9=50\]

OpenStudy (anonymous):

except that is wrong! add 9 divide by 7 sorry

OpenStudy (anonymous):

\[f^{-1}(x)=\frac{x+9}{7}\]

OpenStudy (anonymous):

f(x) = 7x-9 which is the same as y=7x-9 first you always interchange y and x and then you make y the subject of the formula x=7y-9 x+9 = 7y (x+9)/7 = y

OpenStudy (anonymous):

okay. thank you!